OpenStudy (anonymous):

Compute the number of vinyl tiles, measuring 28 inches on each side, needed to tile a kitchen measuring 102 ft. by 18 ft. tiles.

OpenStudy (anonymous):

That is what I need to find out I think.

OpenStudy (alexalas):

so change inches to feet so 28/12=?

OpenStudy (anonymous):

Thats' not a selection.

OpenStudy (alexalas):

I know but it is part of the problem the square is the tile then the area of a square is s^2 one side of the square is 28 inches long we must change it to feet to work with the rest of the equation. so 28/12=(2*1/3) s=(2*1/3) so what is s^2 then the area of the kitchen v*w,(102*18) divide the area of the kitchen by the area of the square tile or in a simple equation (102*18/((28/12)^2) hope this makes sense

OpenStudy (anonymous):

no but it did help me

OpenStudy (alexalas):

did you get the answer?